how do you expand this complex number in a standard complex number format?
(1/2-1/2i)^4

Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(28595)

Instead of calculating (1/2-(1/2)i)^4, I'm going to calculate 4*(1/2-(1/2)i)^4

4*(1/2-(1/2)i)^4

(sqrt(2))^4*(1/2-(1/2)i)^4

(sqrt(2)*(1/2-(1/2)i))^4

(sqrt(2)/2-(sqrt(2)/2)i))^4

(cos(pi/4)-i*sin(pi/4))^4

cos(4*pi/4)-i*sin(4*pi/4)

cos(pi)-i*sin(pi)

-1-i*(0)

-1


So 4*(1/2-(1/2)i)^4 = -1.

But we really want to know the value of (1/2-(1/2)i)^4

So all we need to do is divide both sides by 4 to get

(1/2-(1/2)i)^4 = -1/4


So the final answer is

Answer by solver91311(16897)



Use FOIL to square the binomial just like any other binomial. Remember that . Then square the result.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism